OpenStudy (anonymous):

Write the equation of a trend line. {(-15,8), (-8,7), (-3,0), (0,0), (7,-3)} @hlambach

OpenStudy (hlambach):

Try graphing it first. Plot the points and connect them.

OpenStudy (hlambach):

Naw, don't do that never mind, haha.

OpenStudy (anonymous):

Awww it will take so long! is there any way we can speed it up?

OpenStudy (hlambach):

Yeah, don't do it.

OpenStudy (anonymous):

lol! :D

OpenStudy (hlambach):

This will be a straight line, right?

OpenStudy (anonymous):

Im not sure, will it? :P

OpenStudy (hlambach):

Wait, no, it's not because it is a trend line.

OpenStudy (anonymous):

okay.. not to be rude, but I have to turn this in in 20 mins and ll have 4 questions left. please just this time help me with the answers please this will make or break my grade!

OpenStudy (hlambach):

Okay, so you want to find a slope-intercept equation.

OpenStudy (anonymous):

Are you sure? hey you have 56 not 55 now! :)

OpenStudy (hlambach):

So find the slope of two points using the formula \[(y _{2}-y _{1})/x _{2}-x _{1})\]

OpenStudy (hlambach):

Yay! haha

OpenStudy (anonymous):

could you find the answer while i brb? i have to attend my grandma

OpenStudy (hlambach):

I don't have a calculator with me...

OpenStudy (hlambach):

I have to leave in like ten minutes. After that, just plug in a point in the equation and find b.

OpenStudy (anonymous):

okay i am back.. no you're leaving??? please finish this last one for me!!

OpenStudy (hlambach):

Do what I said and you got it.

OpenStudy (hlambach):

Pick any two points, find the slope, then plug in the slope and a point in the equation y=mx+b

OpenStudy (anonymous):

okay... could you help?

OpenStudy (hlambach):

Yeah. But try it first.

OpenStudy (anonymous):

okay i pick 0,0.

OpenStudy (hlambach):

Now pick another one and find the slope using the formula I gave you earlier.

OpenStudy (anonymous):

okay.. i pick 7, -3 0- 7 is 0 and 0 - -3 is 3

OpenStudy (hlambach):

0-7=-7

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

have 7 apples take away 0 still have 7

OpenStudy (hlambach):

(0-(-3))/(0-7) is how you find teh slope.

OpenStudy (hlambach):

But you put the zero in front.

OpenStudy (anonymous):

-0.428571429 but i have to write an equation of a trend line

OpenStudy (hlambach):

I know, I know. Now see if your calculator can make that into a fraction

OpenStudy (anonymous):

how?

OpenStudy (hlambach):

usually it's a button that looks like this f->d or d->f

OpenStudy (anonymous):

it says the number i entered was too small

OpenStudy (hlambach):

Oh wait, -7/3 is a fraction haha. So that is your slope

OpenStudy (anonymous):

okay how do i find the equation for a trend line? because that is what i need

OpenStudy (hlambach):

Okay, so now that you have a slope for equation, just put in a point in the equation y=-7/3x+b and find b

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

y=-7/3x+(-8,7) now what

OpenStudy (hlambach):

No, put in 7 for y and -8 for x

OpenStudy (hlambach):

In the formula I gave you.

OpenStudy (anonymous):

-8 - 7 is 1

OpenStudy (hlambach):

Is that what b equals? If yes, then you have your formula.

OpenStudy (anonymous):

i think so, am i right?

OpenStudy (anonymous):

y=-7/3x+1?

OpenStudy (hlambach):

yup

OpenStudy (anonymous):

awesome! do you have to go?

OpenStudy (hlambach):

Yeah.

OpenStudy (anonymous):

okay thanks for all the help

OpenStudy (hlambach):

No problem. Thanks for all the medals!

OpenStudy (anonymous):

of course! maybe later you could help more!